# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ 4.4 Experiments and Evaluation
Most approaches train and evaluate using grid2op environments, since they offer extensive power grid simulations with real data. The analyzed works consider various power grid sizes: Xu et al., (2020) focuses on an IEEE 14-bus system with 20 transmission lines, 6 generators, and 11 loads, spanning 1004 scenarios over 4 weeks at 5-minute intervals. Taha et al., (2022), Xu et al., 2022a , and Zhao et al., (2022) employ the larger IEEE 118 grid. Taha et al., (2022) uses the "l2rpn_neurips_2020_track2_small" environment with adjustments, while Sar et al., (2023) uses the IEEE 5 grid for a hierarchical multi-agent proof of concept. Yoon et al., (2021) tackles multiple sizes including IEEE 5, IEEE 14, and an augmented version of IEEE 118 from L2RPN WCCI 2020. Wu et al., 2023b applies their method on IEEE 14 and IEEE 30 bus systems with wind power data for power flow control using Battery Energy Storage Systems (BESS).
In terms of evaluation, the presented approaches on grid control are not as comparable as one could hope, considering that most are based on the same framework and utilize similar data.
@@
Xu et al., (2020) evaluate their approach using a simulation-constraint double dueling DQN agent compared to a basic double dueling DQN agent on an IEEE 14-bus system. The simulation-constraint agent outperforms the basic agent, maintaining grid stability for longer durations per episode. They also assess the effectiveness of graph convolution by training agents with and without GNN layers, demonstrating that the GNN-enabled agent achieves superior control capabilities. While the current results show promise, further assessment on larger-scale grids is necessary to validate the approach’s applicability and effectiveness.
@@
In contrast, Taha et al., (2022) trained on the larger IEEE-118 grid. The GNN input features represent power lines, with a reduced injection horizon for speed improvement. To gauge topology generalization, the GNN is tested on topologies differing by actions from the reference, revealing RMSE evolution logarithmically increasing with action distance, stable up to approximately 7 actions. Although comparison with feedforward neural networks for line load prediction is absent, their interpretable MCTS agent reduces failure rates significantly, from 15.1% to 1.5%, demonstrating the effectiveness of combining GNNs for prediction and MCTS for action optimization, setting the stage for model-based RL using GNNs as model estimators.
@@
Yoon et al., (2021) validated their Semi-Markov Afterstate Actor-Critic (SMAAC) approach against various baselines, including DDQN and SAC, which performed poorly on medium and large grids due to inefficiencies in exploring all possible actions and potential failures. While comparisons between GNN-based and non-GNN RL methods were conducted, detailed validation of specific GNN architectures and alternative models was lacking. Another method, SMAAC/AS, attempted to improve by incorporating goal topology without afterstate representation, but its poor performance underscored the effectiveness of afterstate representation. YZM, another baseline from the same L2RPN challenge, initially showed promise by limiting the considered actions but struggled to further improve due to the challenge of exploring a vast action space. SMAACsignificantly outperformed all other methods, highlighting its effectiveness in learning and managing complex power grids. Additionally, the study explored how different low-level control rules impact overall performance, finding that the high-level SMAAC policy can adapt to various low-level rules as long as they partly achieve goals.
@@
In Sar et al., (2023), two model-free algorithms ( Soft Actor-Critic Discrete (SACD), PPO) are extensively evaluated and compared in their independent and dependent variants within the multi-agent framework. Independent versions have each agent with its own critic, actor, and replay buffer, but struggle with coordination and stability. Dependent versions use a centralized critic to improve coordination among agents, enhancing stability and performance. SACD achieves optimal scores in its single-agent form but is unstable in multi-agent setups, except for DSACD with tuned parameters. PPO, both single- and multi-agent, converges effectively, with single-agent PPO showing faster convergence. Moreover, PPO shows less sensitivity to hyperparameter changes in the multi-agent setting. The model architecture, particularly the GNN employed is not further validated against feedforward networks or other GNN architectures.
@@
Zhao et al., (2022) employs GraphSAGE networks that are evaluated on a modified IEEE 118-bus system. The study includes training these networks unsupervised under specific scenarios and testing them on unseen scenarios with varying grid topologies. They use 2D t-SNE visualization to show the network’s robust representation of the power grid across different setups. Comparing their GraphSAGE-based DRL framework with a dense-based PPO algorithm, they find that GraphSAGE maintains high performance even with changing grid structures, unlike the dense-based approach, which struggles to adapt effectively. However, the evaluation primarily focuses on training outcomes rather than specific power grid performance metrics such as survival time of the agent.
@@
In Xu et al., 2022a , their simulation-driven GRL method is evaluated using the L2RPN Robustness Track challenge dataset. The method shows improved performance over training episodes, combining decisions of sub-agents for collaborative decision-making. Compared to a "do-nothing" approach, it effectively prevents overloads across various scenarios. The model demonstrates rapid decision-making with an average time of 35 ms per time step. They analyze GAT and GCN-based models, finding that GAT supports stable power grid operation better, offering improved stability and economic benefits. However, a comparison with feedforward neural networks is missing. Their Long-Short-Term Action Deployment Strategy outperforms fully reward-guided and enumeration strategies by managing overloads effectively with fewer actions. The chosen action threshold of $0.98$ is validated to yield optimal results.
@@
Wu et al., 2023b evaluated their GRL approach for managing BESS in against baseline methods like DQN and DDPG. They also compare different neural network architectures, focusing on the advantages of their spatio-temporal Chebyshev GCN (Cplx-STGCN) over feedforward, convolutional, and recurrent networks for extracting key features in power grid management. Additionally, the study tests two hybrid OPF solver methods, DeepOPF and DC3. They compare these methods with existing RL approaches like TD3, assessing metrics such as testing rewards and control over power generation and voltage magnitude. Overall, their constrained GRL framework shows superior performance compared to traditional optimization methods and existing RL techniques in managing BESS in power systems.
@@ 4.5 Discussion
The analyzed approaches maintain consistency in state representation aligned with the grid2op environment, yet they vary in the specific information included. For example, Xu et al., (2020) integrates extensive grid features like topology, generator characteristics, and power flow, while others such as Yoon et al., (2021) focus on afterstate representations to capture deterministic changes post-action. This adaptability of GRL methods reflects their ability to handle diverse inputs within the framework.
Generally, the presented approaches heavily rely on the grid2op framework Donnot, (2020) for training and testing, which is designed by transmission grid operators but doesn’t fully model all real-world grid aspects. Consequently, the rewards, actions, states, and graph representations are constrained by grid2op functionalities. Notably, while graph representation significantly affects GNN performance, there’s a lack of analysis on optimizing grid representations within these frameworks.
The diversity of RL methods is evident, with each approach employing tailored algorithms for complex power grid control. For instance, Xu et al., (2020) uses Double-Q-learning with soft constraint enforcement, while Taha et al., (2022) and Xu et al., 2022a employ MCTS for optimal action sequencing. To manage expansive action spaces, techniques like afterstate representations and expert knowledge-based action reduction are utilized to avoid unnecessary interventions in stable grid conditions. However, it is noteworthy that generally there is a limited application of model-based techniques, such as in Fuxjäger et al., (2023). Notably, no model-based technique employing GNNs has been developed so far. Similarly, it is important to highlight that state-of-the-art model-free RL methods, such as BBF Schwarzer et al., (2023), have not been applied.
Exploration poses challenges in environments with strict physical constraints, often leading to violations. Strategies like simulating multiple actions or using surrogate models for faster predictions mitigate these risks. Rewards vary across approaches, reflecting trade-offs in optimizing multiple objectives, with RL enabling comprehensive integration of network control aspects despite potential trade-offs between conflicting objectives.
@@
GNNs are pivotal for extracting features from power grid topologies, enhancing convergence and generalizability across different grid configurations. While GraphSAGE, GATs, and transformer-based models are commonly employed, their evaluation against alternative architectures or feed-forward networks sometimes lacks depth, suggesting opportunities for further comparative studies to ensure robustness and applicability.
@@
Challenges in evaluating GRL approaches include methodological diversity, the need for comprehensive validation, and scalability concerns. Some studies offer thorough evaluations of RL algorithms and architectures, while others focus more narrowly on effectiveness against specific baselines. Additionally, the widely used grid2op environment exhibits significant stochasticity that notably affects agent performance. Lehna et al., (2023) demonstrates performance variation across different random seeds and advocates for averaging results over multiple seeds (e.g., 30) to ensure reliability. However, some approaches discussed overlook this necessity, which could result in incomplete conclusions.
@@
In terms of applicability, it should be noted that the sizes evaluated here are typically smaller than real-world grids, even for the larger test grids. Therefore, the proposed approaches are not yet scalable to real-world grid operations and serve more as a proof of concept. However, they pave the way for GRL-based decision support for grid operators.
@@ 5 Graph Reinforcement Learning for Distribution Grids
Distribution grids, which deliver electricity from the transmission grid to end users, are significantly impacted by the energy system’s transformation. Power generation has increasingly shifted from the transmission system to the distribution side Beinert et al., (2023) due to the rise of distributed renewable energy sources like photovoltaics. This shift causes voltage fluctuations that can threaten grid stability, as system voltages must remain within certain limits for safe operation. New consumers, such as distributed storages and electric vehicles, further exacerbate these fluctuations. Voltage control addresses these issues by flattening voltage profiles and reducing network losses using devices like voltage regulators, switchable capacitors, and controllable batteries Fan et al., (2022). Alternatively, topology control can reduce the need for additional online voltage control devices Xu et al., 2022b . Traditional voltage regulation devices are not designed for highly variable conditions and wear out quickly with frequent use. Therefore, distributed energy resources often include smart inverters that regulate voltages at a fast timescale.
These opportunities encourage the development of novel voltage control methods Srivastava et al., (2023). While DRL has shown promise in this area Duan et al., (2020), the use of GNNs is still emerging. We point out that DRL methods serve as a proof-of-concept and are still very far from practical deployment. Since voltage control involves sequential decision-making to determine optimal voltage set points, DRL is seems promising. Feed-forward neural networks struggle with the topological changes in distribution grids, making the combination of DRL and GNNs a potential better choice to capture the grid’s inherent graph structure.
Table 2: Overview of GRL approaches proposed for voltage-control applications in the distribution grid. The action column lists devices or variables modified by predicted actions (q for reactive power, p for active power). Grid sizes refer to the number of buses evaluated. The focus highlights key aspects or major differences from other approaches.
Model RL Action GNN Grid Size Focus/ Unique Feature
@@
Xing et al., 2023a DDPG p (generators, PV, flexible loads), q (SVC) GAT IEEE 33, 119 consider many objective in reward
Xu et al., 2022b DQN grid topo spectral GCN IEEE 33, 69, 118 action space reduction via GNN + branch exchange
@@ 5.1 Voltage Control
Voltage control is the key task in the distribution grid, managing reactive power set points to maintain grid stability. Reactive power supports voltage levels for delivering real power. The cost function typically includes indicators like power losses and congestion Srivastava et al., (2023). Keeping voltages within limits is crucial to avoid system damage.
The AC power flow in a grid is modeled by highly non-linear equations, making the optimization problem non-convex. To simplify, it is often linearized using methods like semi-definite programming or second-order case programming. For exact models, numerical methods like Newton-Raphson or Gauss-Seidel are used. Heuristics, such as particle swarm optimization, address the non-convexity of the problem formulation Srivastava et al., (2023). Deep learning approaches offer an effective alternative since they suited for non-linear problems, overcoming the limitations of traditional methods in handling the complexity and dynamics of smart grids.
In this work, we distinguish between two cases of voltage control: operation control and emergency mode. In operation control case, i.e. a stable grid state, voltage control is typically addressed via the control of reactive power. In contrast, emergency situations require a more drastic measures such as load shedding, i.e., the cutting of loads to prevent the grid from blacking out.
@@ 5.1.1 Operation Control
Recent studies show that the application of GRL in voltage control for distribution grids. These studies highlight the challenges posed by high levels of renewable energy integration, aiming to minimize network loss, mitigate voltage oscillations, and keep operational costs low. Reactive power control of generators and loads is widely used to manage bus voltage, directly influenced by local reactive power. With future inverter-based generation and digitized grids, maintaining voltage within the desired range (e.g., $0.9$ to $1.1$ p.u.) can be achieved through reactive power control.
@@
Most approaches in this area utilize a similar reward function to encode these objectives. Grid stability hinges on maintaining voltages within specified limits, leading these approaches to penalize voltage deviations from a set reference value. Wu et al., 2023a and Wang et al., (2023) focus solely on these voltage-based rewards. Many methods combine penalties for voltage deviation with other terms in a weighted sum to address multiple aspects supporting grid stability and cost minimization. For instance, Yan et al., (2023) and Mu et al., (2023) incorporate penalties for overall power loss, while Lee et al., 2022b add penalties to discourage frequent control changes that could cause equipment wear. Mu et al., (2023) include a voltage barrier function to constrain voltage ranges.
Meanwhile, Wu et al., 2022c and Li et al., (2023) integrate penalties for PV curtailment and voltage oscillation, with Li et al., (2023) also considering renewable energy integration and generation costs. Cao et al., (2023) base their reward on voltage deviation and use a surrogate model to estimate voltage and power loss. Balancing these reward components is crucial as optimizing one can affect others negatively, often requiring thorough parameter search.
Although Xu et al., 2022b focus on grid stability through changes in power grid topology, they too base rewards on voltage offsets and constraint violations. They introduce a penalty for recommending actions on already disconnected lines and define rewards based on tree metrics to ensure topological actions adhere to certain properties, such as avoiding loops or disconnected
@@
Voltage regulation primarily involves adjusting actuator set points. Approaches proposed by Yan et al., (2023), Mu et al., (2023), Wu et al., 2022c , Cao et al., (2023), Wu et al., 2023a , and Xing et al., 2023a adjust reactive power outputs of PV inverters. Xing et al., 2023a also control the active power of Energy Storage Systems (ESS) and flexible loads, along with reactive power of static var compensation. Cao et al., (2023) additionally predicts set points for static var compressors. Some methods, like Lee et al., 2022b , Wang et al., (2023), and Cao et al., (2023), learn to manage ESS. Lee et al., 2022b also consider modifying configurations of voltage regulators and capacitors.
Alternatively, adjusting active power of generators, including renewables like wind and PV (Li et al., (2023), Wang et al., (2023)), or diesel generators (Wang et al., (2023)) alongside reactive power, is another option. For instance, Li et al., (2023) manipulate active power and voltage magnitude of various generators. In contrast, Xu et al., 2022b focus on modifying grid topology by disconnecting and reconnecting lines.
@@
GRL approaches predict actions based on states such as voltage measurements, load demand, and power generation (both active and reactive). Methods like Lee et al., 2022b , Xu et al., 2022b , and Wu et al., 2023a also consider the status of actuators such as ESS or tap changers. In contrast, Wang et al., (2023) incorporates electricity grid prices. All approaches typically begin by embedding the distribution grid’s current state using a GNN, which encodes feature vectors of states and the grid’s topology represented by the adjacency matrix. These encoded representations are then used by the RL algorithm for decision-making.
@@
The RL algorithms used in these studies primarily include Actor-Critic, PPO, and Q-Learning paradigms, each tailored to specific setups and objectives. For instance, Lee et al., 2022b , Wu et al., 2022c , Xing et al., 2023a , and Li et al., (2023) adopt a straightforward approach using GCNs to embed the distribution grid and train policies. Li et al., (2023) and Xing et al., 2023a employ DDPG for training, while Lee et al., 2022b and Wu et al., 2022c utilize PPO agents. Xing et al., 2023a further emphasize computational efficiency for real-time scheduling.
Apart from single-agent solutions, Yan et al., (2023) and Mu et al., (2023) address zoned distribution networks using a multi-agent Actor-Critic approach, where each agent manages a feeder. Yan et al., (2023) employs a centralized training approach with decentralized decision-making, ensuring physical constraints via a primal-dual method similar to that of Wu et al., 2023b , as described in Sec. 4.2. Similarly, Mu et al., (2023) integrates GCNs into the critic for modeling agent interactions. Wang et al., (2023) employs a multi-agent PPO setups for a distribution grid with multiple microgrids, where each is controlled by an agent that manages its power schedule.
On the other hand, Wu et al., 2023a proposes a two-stage approach: day-ahead optimization using Mixed Integer Second Order Cone Programming, followed by actor-critic learning for voltage regulation in a continuous action space with GCN-enhanced grid embeddings. Cao et al., (2023) also employs a GCN but follow it with a fully connected deep autoencoder for feature dimension reduction in an actor-critic framework, focusing on mapping grid representations to control actions.
In contrast, Xu et al., 2022b manages voltage control by adjusting grid topology. They address the NP-hard complexity, which grows exponentially with grid size, using a method to reduce the action space. They employ deep Q-Learning with a GCN-based Q-network to predict which line to disconnect. For reconnection, a branch exchange based on graph theory is used, adhering to radial constraints that ensure a tree-like network structure without loops. his approach involves selecting lines for reconnection from a basic cut set to maximize rewards, effectively reducing the action complexity.
@@
The GNN architectures in GRL vary significantly across the presented approaches. These GNNs primarily extract features from graphs representing power grids, where nodes correspond to buses and edges to power lines. Nodes include grid information such as voltage, active and reactive power, and the status of controllable actuators (voltage regulators, capacitors, batteries).
For instance, Lee et al., 2022b use a GCN in a straightforward approach with a PPO algorithm and augment their graph with additional edges to expedite information propagation. Similarly, Xing et al., 2023a employ a GAT (cf. Eq. 3) to capture topology branch correlations and power information and trained with DDPG.
Xu et al., 2022b utilize deep Q-Learning with a GCN-based Q-network to predict the best line for disconnection, using spectral convolutions (cf. Eq. 5) for feature extraction. Similarly, Yan et al., (2023) and Wu et al., 2022c apply spectral graph convolutions to merge topology and feature information in Actor-Critic setups. The spectral convolutions act as a low pass filter that suppresses noise in the input data, while the message passing aggregates the features of neighbouring nodes and fills in missing values. Wu et al., 2022c further include a graph shift operator based on AC power flow equations, embedding voltage angle and magnitude. This operator, combined with learnable weight matrices, improves grid feature extraction. They also incorporate temporal information by aggregating node embeddings from the past 10 time steps. Wu et al., 2023a and Li et al., (2023) also introduce novel graph shift operators in their GCNs, with Li et al., (2023) incorporating spatio-temporal attention mechanisms to capture both spatial influences and temporal dependencies across different resolutions.
On the other hand, Mu et al., (2023) employ a tree structure to represent the grid for their multi-agent actor-Critic approach, where a 2-layer spectral GCN integrates information from neighbouring agents. In another multi-agent study, Wang et al., (2023) use a GCN with an actor-critic approach to extract critical network topology and system information for decentralized microgrid voltage regulation. They select critical buses based on their connection to generators, microgrids, or feeder endpoints. These critical buses form the nodes in the grid’s graph. Electrical distance, calculated from the sensitivity coefficients derived from the Jacobian matrix, identifies connections between critical buses. This information is integrated into the adjacency matrix. The GCN, consisting of two graph convolutions, extracts features from the distribution grid, followed by max pooling and two feed-forward layers to learn a control policy.
As an alternative to directly learning the GNN weights using the RL loss, Cao et al., (2023) use a supervised GCN trained on historical power flow data to predict node voltages, which are then used in an actor-critic algorithm for feature extraction without updating the model weights.

### v3

@@ 4.4 Experiments and Evaluation
Most approaches train and evaluate their agents on grid2op because of its extensive power grid simulations with realistic data. Xu et al., (2020) focuses on an IEEE 14-bus system with 20 transmission lines, 6 generators, and 11 loads across 1004 scenarios over 4 weeks at 5-minute intervals. Taha et al., (2022), Xu et al., 2022a , and Zhao et al., (2022) use the larger IEEE 118 grid, while Sar et al., (2023) uses the IEEE 5 grid for a hierarchical multi-agent proof of concept. Yoon et al., (2021) addresses IEEE 5, IEEE 14, and an augmented IEEE 118 grid. Wu et al., 2023b applies their method to IEEE 14 and IEEE 30 bus systems with wind power data for power flow control using Battery Energy Storage Systems (BESS). In terms of evaluation, however, the presented approaches on grid control are not as comparable as one could hope, considering that most are based on the same framework and utilize similar data.
@@
Xu et al., (2020) compare their simulation-constraint double dueling DQN agent with a basic double dueling DQN agent. The simulation-constraint agent outperforms the basic agent, maintaining grid stability for longer durations per episode. They also find that agents with GNN layers outperform those without. While results are promising, further testing on larger grids is needed to confirm the approach’s effectiveness.
@@
Taha et al., (2022) trained their GNN using features representing power lines and a reduced injection horizon for speed. To gauge topology generalization, the GNN was tested on topologies differing by actions from the reference, showing RMSE increases logarithmically with action distance. Although no comparison with feedforward neural networks is provided, their MCTS agent significantly reduces failure rates from 15.1% to 1.5%, demonstrating the effectiveness of combining GNNs for prediction and MCTS for optimization, and paving the way for model-based RL with GNNs.
@@
Yoon et al., (2021) validated their Semi-Markov Afterstate Actor-Critic (SMAAC) approach against baselines like DDQN and SAC, which underperformed on medium and large grids due to inefficient action exploration and potential failures. While they compared GNN-based and non-GNN methods, detailed validation of specific GNN architectures was lacking. The SMAAC/AS method, which incorporated goal topology without afterstate representation, performed poorly, highlighting the value of afterstate representation. Another baseline from the same L2RPN challenge struggled with the vast action space despite initial promise. SMAAC significantly outperformed other methods, proving effective in managing complex power grids and adapting to various low-level control rules.
@@
In Sar et al., (2023), Soft Actor-Critic Discrete (SACD) and PPO are evaluated in both independent and dependent multi-agent settings. Independent agents, each with their own actor, critic, and replay buffer, face coordination and stability issues. Dependent versions use a centralized critic for better coordination, enhancing stability and performance. SACD performs well in a single-agent setting but is unstable in multi-agent scenarios, except for DSACD with tuned parameters. PPO converges effectively in both single- and multi-agent settings, with faster convergence in single-agent and less sensitivity to hyperparameters in multi-agent scenarios. The GNN used is not compared to feedforward networks or other GNN architectures.
@@
Zhao et al., (2022) uses GraphSAGE networks on a modified IEEE 118-bus system, training them unsupervised and testing on various unseen grid topologies. They use 2D t-SNE to demonstrate the network’s robust representation across different setups. Compared to a dense-based PPO algorithm, the GraphSAGE-based method performs well even with changing grid structures, while the dense-based approach struggles to adapt effectively. The evaluation focuses on training outcomes rather than power grid performance metrics like agent survival time.
@@
Xu et al., 2022a evaluate their simulation-driven GRL method using the L2RPN Robustness Track challenge dataset. The method, which combines decisions from sub-agents, prevents overloads more effectively than a "do-nothing" approach and achieves rapid decision-making with an average time of 35 ms per step. GAT models show better stability and economic benefits compared to GCN, though no comparison with feedforward networks is provided. Their Long-Short-Term action deployment strategy outperforms fully reward-guided and enumeration strategies by managing overloads with fewer actions, and the action threshold of $0.98$ is validated as optimal.
@@
Wu et al., 2023b evaluated their GRL approach for managing BESS against baseline methods such as DQN and DDPG. They compare their spatio-temporal Chebyshev GCN (Cplx-STGCN) with feedforward, convolutional, and recurrent networks, highlighting their effectiveness. The study also tests hybrid OPF solvers, DeepOPF, and DC3, comparing them with RL methods like TD3, assessing metrics such as testing rewards and control over power generation and voltage magnitude. Their constrained GRL framework outperforms traditional optimization and existing RL techniques.
@@ 4.5 Discussion
Almost all approaches rely on the Grid2Op framework Donnot, (2020) for training and testing, which, although designed by transmission system operators, abstracts from real-world grid aspects. Therefore, rewards, actions, states, and graph representations are limited to the functionalities provided by Grid2Op. Most studies use similar state representations but differ in the specific information included. For example, Xu et al., (2020) includes extensive features such as topology and power flow, while Yoon et al., (2021) focuses on after-state representations of deterministic changes. This demonstrates the flexibility of GRL to handle different inputs. Despite the importance of graph representations for the performance of GNNs, the studies lack an analysis of how grids are optimally represented as graphs.
@@
GNNs are pivotal for extracting features from power grids, enhancing convergence and generalizability across different configurations. While GraphSAGE, GATs, and transformer-based models are commonly employed, their evaluation against alternative architectures or feed-forward networks sometimes lacks depth, suggesting opportunities for further comparative studies to ensure robustness and applicability.
The diversity in RL methods is clear, with each approach using different algorithms. For example, Xu et al., (2020) uses Double-Q-learning with soft constraints, while Taha et al., (2022) and Xu et al., 2022a employ MCTS for action sequencing. Techniques like afterstate representations and expert knowledge help manage large action spaces and avoid unnecessary interventions. However, model-based techniques, such as those in Fuxjäger et al., (2023), are rarely used, and no model-based approach with GNNs has been developed yet. Additionally, advanced model-free methods, e.g., BBF Schwarzer et al., (2023) have not been applied.
@@
Evaluating GRL approaches faces challenges due to methodological diversity, the need for thorough validation, and scalability issues. Some studies provide in-depth evaluations of RL algorithms and architectures, while others focus on specific baselines. The stochasticity of the Grid2Op environment has a significant impact on agent performance. Therefore, Lehna et al., (2023); Lehna et al., 2024a suggest averaging results over multiple random seeds to increase reliability. However, some approaches neglect this, potentially leading to incomplete conclusions.
@@
In terms of applicability, it should be noted that the evaluated grid sizes are smaller than real-world transmission grids, even for the larger test grids. Therefore, the proposed approaches are not yet scalable to real-world grid operation and serve more as a proof of concept. However, they pave the way for GRL-based decision support for grid operators.
@@ 5 Graph Reinforcement Learning for Distribution Grids
Power generation has increasingly shifted from the transmission system to the distribution side Beinert et al., (2023) due to the rise of distributed renewable energy sources like photovoltaics. This shift causes voltage fluctuations Srivastava et al., (2023) that can threaten grid stability, as system voltages must remain within operational limits. Voltage control addresses these issues by flattening voltage profiles and reducing network losses using devices like voltage regulators, switchable capacitors, and controllable batteries Fan et al., (2022), as well as topology control Xu et al., 2022b . RL is particularly promising for handling multiple objectives in voltage control optimization problems. While DRL has shown promise in this area Duan et al., (2020), the combination with GNNs is still emerging. We point out that DRL methods serve as a proof-of-concept and are still far from practical deployment.
Table 2: Overview of GRL approaches for distribution grids. The action column lists devices or variables modified by predicted actions (q for reactive and p for active power, ESS for energy storage systems, SVC for static var compressors). The column Grid size lists the number of grid buses and Focus/Unique Feature highlights key aspects or major differences to other approaches.
Approach RL Action GNN Grid Size Focus/ Unique Feature
@@
Xing et al., 2023a DDPG p (generators, PV, flexible loads), q (SVC) GAT IEEE 33, 119 Computational efficiency, many objectives in reward
Xu et al., 2022b DQN Grid topology Spectral GCN IEEE 33, 69, 118 Action space reduction via GNN + branch exchange
@@ 5.1 Voltage Control
Voltage control is the key task in the distribution grid, managing reactive power set points to maintain grid stability. While active power is the power that runs devices, reactive power is required to provide the voltage levels that enable the delivery of real power. The cost function in these tasks typically includes system-wide indicators such as power losses and congestion Srivastava et al., (2023). It is essential that the voltages remain within the prescribed limits, as any violation would have detrimental effects on the system.
The AC power flow in a grid is modeled by highly non-linear equations, making the optimization problem non-convex. To simplify, it is often linearized using methods like DC-OPF formulations Tinney and Hart, (1967). For exact models, numerical methods such as Newton-Raphson or Gauss-Seidel are used. Heuristics, such as particle swarm optimization, address the non-convexity of the problem formulation Srivastava et al., (2023). Deep learning approaches offer an effective alternative since they are suited for non-linear problems, overcoming the limitations of traditional methods in handling the complexity and dynamics of smart grids.
In this work, we distinguish between two cases of voltage control: operation control and emergency mode. In the operational control case, i.e. a stable grid state, voltage control is typically addressed via the control of reactive power. In contrast, emergency situations require more drastic measures such as load shedding, i.e., the cutting of loads to prevent the grid from blackout. Tab. 2 lists the RL method, action type, GNN architecture, grid size, and overall focus of the analyzed approaches.
@@ 5.1.1 Operation Control
Reactive power control of generators and loads is commonly used to manage bus voltage, influenced by local reactive power. Future inverter-based generation and digitized grids can maintain voltage within the desired range (e.g., $0.9$ to $1.1$ p.u.) through reactive power control, aiming to minimize network loss, mitigate voltage oscillations, and reduce operational costs.
@@
Grid stability relies heavily on maintaining voltages within defined limits, so all approaches penalize voltage deviations from a set reference value. Many methods combine penalties for voltage deviation with other terms in a weighted sum to address multiple aspects that support grid stability and cost minimization. Additional penalties address power loss Yan et al., (2023); Mu et al., (2023), equipment wear Lee et al., 2022b , and voltage barrier functions to constrain voltage ranges Mu et al., (2023). Terms for PV curtailment, voltage oscillation Wu et al., 2022c ; Li et al., (2023), renewable integration, and generation costs Li et al., (2023) are also considered. Cao et al., (2023) base their reward on voltage deviation and use a surrogate model to estimate voltage and power loss. Finally, Xu et al., 2022b focus on grid stability through power grid topology changes, basing rewards on voltage offsets and constraint violations. They penalize actions on already disconnected lines and define rewards based on tree metrics to avoid loops or disconnections. Balancing multiple reward components is critical, as optimizing one can negatively impact others.
@@
Voltage regulation involves adjusting actuator set points. Approaches like Yan et al., (2023); Mu et al., (2023); Wu et al., 2022c ; Cao et al., (2023); Wu et al., 2023a ; Xing et al., 2023a adjust reactive power outputs of PV inverters. while others also control the active power of Energy Storage Systems (ESS) Xing et al., 2023a ; Lee et al., 2022b ; Wang et al., (2023); Cao et al., (2023), flexible loads Xing et al., 2023a and static var compensation Xing et al., 2023a ; Cao et al., (2023). Adjusting the active power of generators of renewables (Li et al., (2023), Wang et al., (2023)) or diesel generators (Wang et al., (2023)) alongside reactive power is another option. In contrast, Xu et al., 2022b focuses on modifying grid topology by disconnecting and reconnecting lines.
@@
The presented GRL methods predict actions based on states such as voltage measurements, load demand, and power generation. The status of actuators, e.g., ESS, tap changers Lee et al., 2022b ; Xu et al., 2022b ; Wu et al., 2023a , and electricity grid prices Wang et al., (2023) are also considered. States are typically embedded using GNNs to encode the distribution grid’s feature and topology. These encoded representations are then used by the RL algorithm for decision-making.
@@
Common RL algorithms in these studies include Actor-Critic, PPO, and Q-Learning each tailored to specific setups and objectives. Approaches like Lee et al., 2022b ; Wu et al., 2022c ; Xing et al., 2023a ; Li et al., (2023) use a GCN for grid embedding and train policies with DDPG or PPO. Multi-agent actor-critic setups Yan et al., (2023); Mu et al., (2023) with one agent per zone manage zoned networks with centralized training and decentralized execution, i.e. based on only local observations. Yan et al., (2023) integrate the GNN into the actor networks, while Mu et al., (2023) use the GCN only in the critic to model agent interactions. To ensure that the physical equations of the physical system are satisfied, a primal-dual method similar to that of Wu et al., 2023b , as described in Sec. 4.2, is used. Similarly, Wang et al., (2023) use multi-agent PPO with a GNN in both actor and critic for microgrid management, where each microgrid is controlled by an agent that manages its power schedule.
Wu et al., 2023a propose a very different two-stage approach: day-ahead optimization using Mixed Integer Second Order Cone Programming, followed by actor-critic learning for voltage regulation in a continuous action space with GCN-based grid embeddings. Cao et al., (2023) also use a GCN to embed the grid, but they apply a subsequent fully connected deep autoencoder for feature dimension reduction in an actor-critic framework. In contrast, Xu et al., 2022b manages voltage control by adjusting grid topology, addressing the NP-hard complexity with a method to reduce the action space. They use deep Q-learning to predict line disconnections and develop a branch exchange mechanism that considers the radial constraints when selecting a line to be reconnected.
@@
The GNN architectures extract features from power grid graphs, where nodes represent buses and edges represent power lines. However, they vary significantly across the presented approaches. Multiple approaches employ the popular GAT (cf. Eq. 3) or some variant of it to capture topology branch correlations and power information Xing et al., 2023a ; Lee et al., 2022b ; Li et al., (2023); Xing et al., 2023a .
Some studies Lee et al., 2022b ; Wang et al., (2023) extend their graph before applying a GNN such as Lee et al., 2022b , which introduce additional edges to allow faster information propagation through the graph. For their decentralized microgrid voltage control, Wang et al., (2023) model their graph so that only critical buses (those connected to generators, microgrids, or feeder endpoints) serve as nodes, connected by edges with edge weights based on electrical distance information.
Several approaches apply spectral graph convolutions (cf. Eq. 6) for feature extraction Xu et al., 2022b ; Yan et al., (2023); Wu et al., 2022c ; Mu et al., (2023). The spectral convolutions act as a low pass filter that suppresses noise in the input data, while the message passing aggregates the features of neighboring nodes and fills in missing values. One multi-agent study uses a tree structure to represent the grid, where a spectral GCN integrates information from neighboring agents Mu et al., (2023).
Contrary to the other approaches, the GNN of Xu et al., 2022b outputs edge instead of node embeddings by aggregating the nodes connected to the edge since the approach controls voltages by disconnecting lines. Wu et al., 2022c propose a novel graph shift operator based on the AC power flow equations, embedding the voltage angle and magnitude. Furthermore, the approach incorporates temporal information by aggregating the node embeddings of the last 10 time steps. Similarly, Li et al., (2023) present a spatio-temporal attention mechanism that learns the temporal dependency between the same node embedding at different time steps. The spatio-temporal attention convolutions are applied sequentially so that the result of the spatial is convolved in the temporal dimension.
Another graph shift operator is introduced by Wu et al., 2023a , but in contrast to Wu et al., 2022c , it is based on the grid topology and the correlation coefficient matrix obtained from the PV and load historical data.
In terms of training, Cao et al., (2023) present an alternative to directly learning the GNN weights using the RL loss. They first train a GCN in a supervised manner on historical power flow data to predict node voltages. Then, the weights of this surrogate model are copied to the representation networks of the actor-critic algorithms to perform feature extraction from the distribution network.